Derek sold walking sticks. Had he charged $10 extra for each stick he would have made $1,800. However, had he sold 10 more sticks at the original price he would have made $2,000.
a) How many sticks did he sell?
b) How much was each stick?

Found 2 solutions by hk, mananth:
Answer by hk(4)

let us assume selling price of walking stick is 'x'and that he sold 'n'no of sticks.
from information given:
(x+10)*n=1800 ----- (I) and
x*(n+10)= 2000 --------(II).
from I, xn+10n = 1800; and from II xn+10x = 2000
solving the equations we get x = n+20. Putting this back in equation I, and solving for n, we get n= 30 (no of sticks sold). Hence price of stick ie. x is 50

Answer by mananth(16946)

Derek sold walking sticks. Had he charged $10 extra for each stick he would have made $1,800. However, had he sold 10 more sticks at the original price he would have made $2,000.
a) How many sticks did he sell?
b) How much was each stick?
let number of walking sticks sold be x
and price of walking stick =y
(y+10)*x=1800
(x+10)*y = 2000
x=1800/y
((1800/(y+10)) +10)*y = 2000
(1800+10y+100)*y/(y+10)=2000
1800y+10y^2+100y=2000y+20000
10y^2-100y-20000=0
/10
y^2-10y-2000=0
y^2-50y+40y-2000=0
y(y-50)+40(y-50)=0
(y-50)(y+40) = 0
y= 50 the positive value the price of the stick
plug y to get x
x= 30 number of sticks